OpenStudy (anonymous):

P = x + 2y + 3z solve for z

OpenStudy (igreen):

Okay, to solve for \(z\), we have to get it by itself. \(P = x + 2y + 3z\) We can subtract \(x\) to both sides: \(P - x = 2y + 3z\) Now we can subtract \(2y\) to both sides: \(P - x - 2y = 3z\) Do you know what we do next?

OpenStudy (anonymous):

not really ^^' i hardly understand the question.

OpenStudy (anonymous):

divide 2y? @iGreen ?

OpenStudy (igreen):

You divide 3 to all the terms to the left. So what does that leave you with?

OpenStudy (anonymous):

p - x - 2y -------- = 3x 3 ??

OpenStudy (anonymous):

*not 3x just x, sorry.

OpenStudy (anonymous):

@iGreen ?

OpenStudy (anonymous):

anyone??

OpenStudy (texaschic101):

P = x + 2y + 3z .... solving for z--- subtract x from both sides P - x = 2y + 3z --- subtract 2y from both sides P - x - 2y = 3z --- now divide both sides by 3 (P - x - 2y) / 3 = z OR P/3 - 1/3x - 2y/3 = z

OpenStudy (texaschic101):

or P/3 - 1/3x - 2/3y = z

OpenStudy (igreen):

Yeah, that's correct. @silverr

OpenStudy (anonymous):

yayy thank you so much you guys !